OpenStudy (anonymous):

hi guys, I am just learning algebra could someone please show me the instructions to find a missing numerator? 3/7 = ?/49 thanks in advance

OpenStudy (anonymous):

You have to get the denominators to equal. They both need to be 49. So what times 7 = 49. Then what ever you do to the numerator you must do to the denominator.

OpenStudy (anonymous):

\[3/7=x/49\] So what this says, it's that the number x divided by 49 is equal to 3/7 So if you multiply both sides(terms) by 49 would result in geting the number x in one side and the equivalent in another \[(3/7)*49=(x/49)*49\] Would give you the number x \[(3/7)*49=x\]

OpenStudy (anonymous):

Think of it like this \[\frac{ 3 }{ 7 } \times \frac{ ? }{ ? } = \frac{ x }{ 49 }\]

OpenStudy (anonymous):

thanks guys brain fried will write both down and look in the morning

OpenStudy (anonymous):

I know 7 * 7 =49

OpenStudy (anonymous):

Any number divided by the same number will give you 1 \[50/50=1\] \[3450/3450=1\] You can move around operations to fit your taste(as long as you don't modify the overall result) So \[(49/49)*x=x\] Just complementing my answer above :)